& If & & w(t)=6t+31, & solve w(t-5) = w(2t). & t = ?
1
Expert's answer
2011-09-07T05:05:52-0400
w(t)=6t+31,
w(t-5) = 6(t-5)+31 = 6t - 30 + 31 = 6t +1
w(2t) = 6(2t) + 31 = 12t + 31
Hence:
6t +1 = 12t + 31
6t = - 30
t = -5